OpenStudy (anonymous):

A circle has its center at the origin, and (5, -12) is a point on the circle. How long is the radius of the circle?

OpenStudy (asadkarim7):

apply distance formula sqrt(x2-x1)^2 + ( y2-y1)^2

OpenStudy (anonymous):

Circle is x^2 + y^2 = r^2 so 25 +144 = r^2, r = 13.